OpenStudy (anonymous):

Hey I have a question, if say you get something like 3x, if x = 4 what would the 3x be? Like 3 to the fourth power, I don't know, HEEEELP D:

OpenStudy (anonymous):

3*4

OpenStudy (anonymous):

is 12

OpenStudy (anonymous):

3x4 3 times 4 so yes it is 12

OpenStudy (anonymous):

good job logan

OpenStudy (anonymous):

Thanks. I learned that last year :P

OpenStudy (anonymous):

THANK YOU GOD THANK YOU :DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

its cause the question was

OpenStudy (anonymous):

What is the value of the expression 3x + x – 10 when x = 4? 58 19 11 6

OpenStudy (anonymous):

so it would be 12+4-10 12+4=16 16-10=6

OpenStudy (anonymous):

i love you.

OpenStudy (anonymous):

lol you too :D

OpenStudy (anonymous):

i fan hardcore ;-;